HI WORKING ON ORDERED PAIRS HERE. THE PROBLEM IT:
Y=7-4X AND I NEED TO FIND 3 DIFFERENT ORDERED PAIRS THAT ARE SOLUTIONS OF THE EQUATION.

Answer by drk(1908)

Let x = -1, 0, and then 1.
If x = -1, we get y = 7 - 4*(-1) = 11
(-1,11)
If x = 0, we get y = 7 - 4*(0) = 7
(0,7)
If x = 1, we get y = 7 - 4*(1) = 3
(1,3)